# Modeling Extensive Defects in Metals through Active Machine Learning and Automated Configuration Reconstruction

### v1

@@ Abstract
Extended defects such as dislocation networks and complex grain boundaries are ubiquitous in metals, and accurately modeling these extensive defects is crucial for understanding their deformation mechanisms. Existing machine learning interatomic potentials (MLIPs) often fall short in adequately describing these defects, as their significant characteristic sizes exceed the computational limits of first-principles calculations. To address these challenges, this study constructs a comprehensive defect genome, derived through active machine learning from classical empirical potentials, and augmented by conventional on-the-fly active learning. Additionally, we have developed a approach for transforming defect atomic clusters into periodic configurations via precise atom insertion, using Grand Canonical Monte Carlo simulations. These strategies enable the development of highly accurate and transferable MLIPs for modeling extensive defects in metals. Using body-centered cubic tungsten as a model system, we develop an MLIP that reveals previously undocumented plastic mechanisms in simulations of nanoindentation and dislocation-grain boundary interactions. This framework not only improves the modeling accuracy of extensive defects in crystalline materials but also establishes a robust foundation for further advancement of MLIP development through the strategic use of defect genomes.
@@ INTRODUCTION
In the intersecting realms of computational chemistry, materials science, and mechanics, machine learning has made substantial strides, particularly through the creation of machine learning interatomic potentials (MLIPs). These innovative tools have transformed atomic-scale simulations by facilitating the accurate modeling of complex materials behavior with quantum-level accuracy  (?, ?, ?, ?, ?, ?). A diverse array of MLIPs leveraging unique descriptors has emerged. For instance, Neural Network Potentials (NNPs) employ atom-centered symmetry functions (ACSF)  (?), while Gaussian Approximation Potentials (GAP) utilize the smooth overlap of atomic positions (SOAP)  (?), and Spectral Neighbor Analysis Potentials (SNAP) depend on bispectrum components  (?). Additionally, both Moment Tensor Potential (MTP)  (?) and Atomic Cluster Expansion (ACE)  (?) have found extensive application across various materials, from metals and alloys to 2D materials and complex systems like carbon and silicon  (?, ?, ?). A Pareto analysis reveals that while GAP achieves the highest accuracy, it lacks computational efficiency  (?). In contrast, MTP and ACE present a more balanced profile of accuracy and computational demand, with ACE being notably faster and less resource-intensive than MTP  (?, ?, ?).
@@
An essential element in developing MLIPs is constructing a good training database that adequately covers the configuration space relevant to the intended simulations. Typically, this database is curated by domain knowledge, such as ground-state structures, structures deformed under various elastic strains, ab initio molecular dynamics (AIMD) configurations at different temperatures, and defects such as vacancy, interstitial, crack tip and dislocation core  (?). Advancements in machine learning have led to the adoption of active learning (AL)  (?, ?, ?), a more sophisticated approach that selects new configurations efficiently by assessing uncertainty without the need for density functional theory (DFT) calculations. The effectiveness of AL strategies hinges on the ability to calculate similarities among configurations by comparing their (local atomic environments) LAEs. MTP with AL, utilizing the D-optimality criterion and the MaxVol algorithm, exemplifies one of the most widely employed MLIPs, which has been used in numerous studies  (?, ?). Other AL strategies, such as uncertainty-driven dynamics  (?) and hyperactive learning  (?), also significantly contribute to efficiently sampling the configuration space.
@@
Two primary challenges persist in the construction of databases for the development of MLIPs. The first challenge is the comprehensive coverage of defect. For metals, it is relatively straightforward to train MLIPs that can reproduce lattice constants, elastic constants, general stacking fault energy, simple grain boundaries, and dislocation migration. However, it remains unclear whether existing MLIPs can effectively model extended defects and their interactions, even if some of them are deemed for general-purpose  (?, ?, ?, ?, ?). This uncertainty primarily stems from the lack of quantification for unknown defects. Although on-the-fly AL (OTF-AL) has been proposed to accelerate the development of MLIPs, it is insufficient to sample comprehensive defects in every loading condition. Qi et al. suggested that an ideal strategy would be to efficiently generate and sample configuration space before conducting any DFT calculations  (?). They used M3GNet  (?) as the engine for effective and efficient database generation for atomic hydrogen diffusion in titanium hydrides system. This strategy however is unsuitable for modeling extended defects in metals due to the low accuracy and computational efficiency of universal MLIPs in large-scale simulations.
@@
In this study, we introduce a generalized framework for developing MLIPs specifically designed to model extensive defects in metals. This framework integrates two key components: (1) Empirical Interatomic Potentials-assisted Active Learning (EIP-AL), and (2) Periodic Configuration Construction via Grand Canonical Monte Carlo simulations (PCC-GCMC). These components are crucial for overcoming the challenges previously discussed. The EIP-AL method is designed to enhance defect sampling capabilities during large-scale simulations across a variety of loading conditions. When combined with conventional OTF-AL, this broad sampling is crucial for capturing the full spectrum of defects and their complex interactions, ultimately facilitating the construction of a comprehensive defect genome. On the other hand, the PCC-GCMC technique converts non-periodic atomic clusters of selected defects, into periodic configurations without vacuum. This conversion is essential for ensuring that these configurations are compatible with standard plane-wave DFT calculations. Focusing on body-centered cubic tungsten (BCC W), our analysis underscores the critical role of the EIP-AL and PCC-GCMC techniques in enhancing the predictive accuracy and reliability of MLIPs for large-scale simulations. Ultimately, we have developed a set of MLIPs optimized for BCC W, capable of simulating a wide range of plastic deformations and crack propagation. We demonstrate the capabilities of our MLIPs through nanoindentation simulations on single crystals and compressive tests on spherical bicrystals featuring both low-angle and high-angle twist GBs. These tests reveal new deformation mechanisms not previously documented, underscoring the robustness and adaptability of our MLIPs in handling complex simulation scenarios.
@@ RESULTS
In this study, we introduce an generalized framework designed to develop versatile machine learning interatomic potentials (MLIPs) for accurately modeling extensive defects in metals, as depicted in Fig. 1. This framework is implemented through a sequential three-step process. Initially, a comprehensive pool of defect structures is generated using large-scale simulations based on well-established EIPs, notably the Embedded-Atom Method (EAM) and the Modified Embedded-Atom Method (MEAM) for metals. These potentials have been instrumental for over three decades, enabling the rapid simulation of complex defects across millions of atoms, albeit with lesser accuracy compared to DFT. Such capabilities are crucial for simulating a variety of deformation scenarios, including but not limited to general GBs in polycrystals, as well as compression, tension, and shear tests in both polycrystalline and single-crystal systems, and nanoindentation (Fig. 1a). This methodological versatility fosters the creation of a diverse set of local atomic environments for defects (LAEDs), addressing phenomena such as grain boundary relaxation and deformation, dislocation nucleation from material inside and exposed surfaces, dislocation multiplication, and complex dislocation-grain boundary interactions. Subsequently, the D-optimality criterion-base AL implemented in MLIP-3  (?) is used to select representative LAEDs based on the local atomic neighborhoods. This strategy meticulously evaluates the extensive pool of generated configurations against an established basic dataset, enabling the precise selection of the most representative LAEDs. This process yields distinct atomic clusters with typical size of 100-200 atoms. For these clusters, vacuum layers of 8 Å thickness are added in all three dimensions of the simulation box to minimize image interaction. We develop an MLIP, named C-MLIP-0, combining the basic dataset with these cluster configurations. This MLIP is also the result of employing the MLIP-3 approach  (?).
@@
In the second step, our PCC-GCMC strategy is employed to convert all clusters from the first step into periodic configurations as shown in Fig. 1b. For each atomic cluster, we begin by initializing the simulation box size slightly larger than the atomic coordinates in all directions. Subsequently, new atoms (depicted in blue in Fig. 1b) are inserted into the box via GCMC simulations, continuing until no additional atoms can be accommodated. During this insertion phase, the core atoms (shown in red) remain fixed to preserve the original LAE of the center atom. After the insertion is complete, both the coordinates of the newly inserted blue atoms and the box size are relaxed simultaneously. This step is crucial to prevent highly distorted LAEs among the newly inserted atoms. This procedure is iteratively repeated, gradually increasing the box size to facilitate further atom insertions. Through this iterative process, the configuration that exhibits the lowest potential energy throughout all iterations is selected as the final configuration for further single-point DFT calculations. The resulting new DFT dataset, combined with the basic dataset, is utilized to train a foundational model, termed P-MLIP-0, designed to be suitable for simulating common defects. To demonstrate the functionality of our approach, we use the example of a screw dislocation core in BCC W to detail the PCC-GCMC, as described in Supplementary Note 1 and illustrated in Fig. S1. Remarkably, without any manual adjustments, our method successfully generates two distinct periodic configurations: one featuring a screw dislocation dipole and another with a non-compact core in addition to the original compact core. These two periodic configurations are readily prepared for DFT calculations and are instrumental for the development of MLIPs specifically tailored for modeling screw dislocation phenomena.
@@
The final stage, Step 3, is the MLIP refinement phase, as illustrated in Fig. 1c. For specific applications not addressed in the initial EIP-AL of Step 1, such as crack propagation and radiation damage, further EIP-AL and PCC-GCMC operations are required, utilizing appropriate EIPs. A limitation of EIP-AL is its potential bias in generating necessary LAEDs due to the inherent bias of EIPs. Consequently, new LAEDs may be generated during MLIP-based large-scale simulations. To ensure the transferability of MLIPs across diverse applications, we implement the traditional OTF-AL process here for specific cases. All generated atomic clusters are converted into periodic configurations via PCC-GCMC and reincorporated into the training dataset. This process is repeated until no new LAEDs are produced by OTF-AL. The final version of our MLIP, named P-MLIP-1, represents the culmination of our development efforts and is designed to robustly model a wide range of defect phenomena.
Figure 1: The framework of versatile MLIP development for modeling extensive defects. a Empirical interatomic potential-informed active learning (EIP-AL). b Construction of periodic configurations from non-periodic atomic clusters through precise atom insertion using grand canonical Monte Carlo simulations (PCC-GCMC). c MLIP refinement by combining EIP-AL, on-the-fly (OTF-AL), and PCC-GCMC.
Structural analysis of initial defect pool based on empirical interatomic potential-based active learning
@@
This section explores the differences between the cluster dataset shown in Fig. 1a and the periodic dataset from Fig. 1b, alongside the basic dataset informed by domain knowledge. We will examine their localized atomic environments (LAEs) and assess MLIP training performance to identify key distinctions. This examination is conducted using the SOAP descriptor  (?) with parameters set as $r_{\text{cut}}$ = 5 Å, $n_{\text{max}}$ = 12, and $l_{\text{max}}$ = 10  (?). The results are visualized in Fig. 2a, where the first two principal components (PCs) are plotted against each other. In the basic dataset, the majority of atomic environments exhibit PC1 values less than 0, with some data points extending into a long arm, representing surface atoms. The cluster dataset, conversely, divides into two groups: a smaller group that aligns closely with the main portion of the basic dataset and a larger group that positions itself above the extended arm of the basic dataset. This larger group primarily comprises surface atoms, a logical outcome given their proximity to the free surface. The smaller group represents target atoms situated at the center of each cluster configuration. The size disparity between these groups indicates that only a small fraction of LAEs directly contribute to target defects, while the majority, being surface atoms, do not inform the force and energy calculations of non-surface atoms. Unlike the cluster dataset, the periodic dataset forms a single cohesive group containing all target atoms from the cluster dataset. There is an overlapping region between the periodic and basic datasets, marked by a dashed line, suggesting shared common LAEs. This overlap signifies that the periodic configurations created via our PCC-GCMC method exhibit characteristics similar to those identified in Ref  (?), with data points in this region likely corresponding to the perfect BCC lattice. Additionally, other data points in the periodic dataset represent inserted atoms from the GCMC process. Although these LAEDs are not explicitly defined in the initial EIP-AL (such as non-compact screw core discovered in Fig. S1), they may prove instrumental in exploring other types of unknown plastic deformations, which we will discuss further in the following section.
@@
Next, we proceed to train two sets of MLIPs, i.e. C-MLIP-0 and P-MLIP-0, utilizing the cluster dataset (C) and the periodic dataset (P), respectively, each integrated with the basic dataset (B). MTP is used for training. The training performance for these models is assessed by comparing the root mean square error (RMSE) of configuration energies, as presented in Fig. 2(b). It is consistently observed that the B+P combination outperforms B+C across various MTP levels. This superior performance of B+P can be attributed to the more straightforward nature of the dataset, which is devoid of the complex, irrelevant LAEs like surface atoms found in the cluster dataset (illustrated in Fig. 2a). This finding underscores the first significant advantage of employing the periodic dataset in MLIP training: the elimination of irrelevant LAEs enhances the training process, leading to more accurate and reliable MLIPs.
Figure 2: Structural analysis of defect pool from EIP-AL. a Principal component analysis of different datasets. b MTP training performance of two datasets with respect to different levels.
To further elucidate the superiority of PCC-GCMC and the resultant P-MLIP-0, we undertake additional validation tasks by comparing the energy and forces predicted by our MLIP-0 models against new DFT calculations for various GBs relaxed at temperatures of 300 K, 900 K, and 2000 K, as well as under severe compression conditions. The stable structures of these GBs are referenced from an established study in Ref  (?). The GB structures and comparative results for energy and force between MLIPs and DFT are displayed in Fig. S2. Observations reveal that both C-MLIP-0 (trained by B+C) and P-MLIP-0 (trained by B+P) consistently outperform B-MLIP-0 (trained solely by B) across all validation scenarios. An exception is noted in the energy comparison at 2000 K, where B-MLIP-0 records a lower error than C-MLIP-0 and P-MLIP-0 (6.3 vs 12.2 and 7.2 meV/atom). However, the force RMSE for B-MLIP-0 is significantly higher compared to C-MLIP-0 and P-MLIP-0 (199.3 vs 172.2 and 154.4 meV/Å). These findings suggest that incorporating either cluster (C) or periodic (P) datasets enhances the accuracy. Notably, P-MLIP-0 consistently surpasses C-MLIP-0 in both energy and force accuracy, particularly under compression conditions (6.7 vs 10.1 meV/atom for energy; 150.8 vs 173.7 meV/Å for force). This superior performance can be attributed to the inclusion of additional atoms by PCC-GCMC in Fig. 2a, which introduces relevant LAEDs absent in B+C, and thus better aligns with the new validation tasks in Fig. S2. Combining the training performance illustrated in Fig. 2b, we conclude that B+P offers superior training outcomes and higher transferability than B+C, thereby validating its enhanced efficacy in modeling complex defects.
Uncertainty quantification in general plastic deformation based on the initial defect pool
@@
Our MLIP-0 models, C-MLIP-0 and P-MLIP-0, are designed to accommodate a wide range of defect simulations including dislocations, GBs, dislocation nucleation and multiplication, and dislocation-GB interactions. To verify the transferability of our MLIP-0 models, we utilize per-atom uncertainty quantification available in MLIP-3  (?), assessing the extrapolation grade, $\gamma$, to evaluate various defects in BCC W. Values of $\gamma$ ranging from 0 to 1 signify interpolation, whereas values exceeding 1 imply extrapolation. Figs 3a-c display three common types of dislocation loops encountered in the plastic deformation of BCC metals: a $<100>$ interstitial loop, a $<111>$ glide loop, and a $<111>$ climb loop. Given that the performance outcomes of the basic dataset combined with cluster one (B+C) and the basic dataset with periodic one (B+P) are comparable, we present only the results for B+C for clarity. Notably, using the basic dataset (B), all atoms within the $<100>$ interstitial loop and those near the edge segments of the $<111>$ glide loop, along with some atoms in the $<111>$ climb loop, exhibit $\gamma$ values significantly greater than 1, indicating a higher level of uncertainty (left panels of Fig. 3a-c). In contrast, all atoms within these dislocation loops, when analyzed using the B+C dataset, show $\gamma$ values less than 1. This demonstrates our MLIP-0 models’ enhanced effectiveness in capturing complex dislocations, with energies and forces of all LAEs being obtainable through interpolation.
@@
Additionally, we assess the transferability of our models regarding general GBs in polycrystals and their temperature-induced evolution. We calculate the extrapolation grade $\gamma$ for all the atoms in a polycrystal relaxed by a MEAM potential  (?) at 300 K and 2000 K. It should be noted that our EIP-AL in the initial EIP-AL only use EAM-Zhou potential  (?), such that the GBs obtained by MEAM may have different atomic structures. Fig. 3d illustrates the polycrystal, with each atom color-coded according to $\gamma$ values obtained from B+C. Notably, all atoms in this representation exhibit $\gamma$ values less than 1, indicating low uncertainty and high confidence in the predictions. Figs 3e and f display histogram plots of the $\gamma$ values for this polycrystal calculated using different datasets: the basic dataset (B), a dataset from a Gaussian Approximation Potential (GAP)  (?), B+C, and B+P at 300 K and 2000 K. The analysis reveals that both datasets B and GAP predict $\gamma$ values greater than 1 for atoms at both 300 K and 2000 K, indicating higher uncertainty. This is particularly pronounced for B. In contrast, $\gamma$ values calculated by datasets B+C and B+P are consistently below 1 at 300 K, suggesting a higher accuracy in these configurations. However, at 2000 K, while most atoms still exhibit $\gamma$ values below 2, there are a few exceptions with values slightly higher than 2. This indicates that although the models perform well at lower temperatures, at higher temperatures the uncertainty escalates but stays manageable. These results underscore the high transferability and effectiveness of generated defect pool in simulating the behavior of general GB and the thermodynamics of polycrystals.

### v2

# Modeling Extensive Defects in Metals through Classical Potential-Guided Sampling and Automated Configuration Reconstruction

@@ Abstract
Extended defects such as dislocation networks and general grain boundaries are ubiquitous in metals, and accurately modeling these extensive defects is crucial for understanding their deformation mechanisms. Existing machine learning interatomic potentials (MLIPs) often fall short in adequately describing these defects, as their significant characteristic sizes exceed the computational limits of first-principles calculations. In this study, we address these challenges by establishing a comprehensive defect genome through empirical interatomic potential-guided sampling. To further enable accurate first-principles calculations on this defect genome, we have developed an automated configuration reconstruction technique. This method transforms defect atomic clusters into periodic configurations through precise atom insertion, utilizing Grand Canonical Monte Carlo simulations. These strategies enable the development of highly accurate and transferable MLIPs for modeling extensive defects in metals. Using body-centered cubic tungsten as a model system, we develop an MLIP that reveals unique plastic mechanisms in simulations of nanoindentation. This framework not only improves the modeling accuracy of extensive defects in crystalline materials but also establishes a robust foundation for further advancement of MLIP development through the strategic use of defect genomes.
@@ INTRODUCTION
In the intersecting realms of computational chemistry, materials science, and mechanics, machine learning has made substantial strides, particularly through the development of machine learning interatomic potentials (MLIPs). These innovative tools have transformed atomic-scale simulations by facilitating the accurate modeling of complex materials behavior with quantum-level accuracy  (?, ?, ?, ?, ?, ?). A diverse array of MLIPs leveraging unique descriptors has emerged. For instance, Neural Network Potentials (NNPs) employ atom-centered symmetry functions (ACSF)  (?), while Gaussian Approximation Potentials (GAP) utilize the smooth overlap of atomic positions (SOAP)  (?), and Spectral Neighbor Analysis Potentials (SNAP) depend on bispectrum components  (?). Additionally, both Moment Tensor Potential (MTP)  (?) and Atomic Cluster Expansion (ACE)  (?) have found extensive applications across various materials, from metals and alloys to 2D materials and complex systems like carbon and silicon  (?, ?, ?). A Pareto analysis reveals that while GAP achieves the highest accuracy, it lacks computational efficiency  (?). In contrast, MTP and ACE present a more balanced profile of accuracy and computational demand, with ACE being notably faster than MTP  (?, ?, ?).
@@
An essential element in developing MLIPs is constructing a good training database that adequately covers the configuration space relevant to the intended simulations. Typically, this database is curated by domain knowledge, such as ground-state structures, structures deformed under various elastic strains, ab initio molecular dynamics (AIMD) configurations at different temperatures, and defects such as vacancy, interstitial, crack tip and dislocation core  (?). Advancements in machine learning have driven the adoption of on-the-fly active learning (OTF-AL)  (?, ?, ?), a sophisticated approach that accelerates the development of MLIPs by efficiently selecting new configurations exclusively for necessary DFT calculations. The effectiveness of OTF-AL strategies hinges on the ability to calculate similarities among configurations by comparing their local atomic environments (LAEs). MTP with OTF-AL, utilizing the D-optimality criterion and the MaxVol algorithm, exemplifies one of the most widely employed MLIPs, which has been used in numerous studies  (?, ?). Other active learning strategies, such as uncertainty-driven dynamics  (?) and hyperactive learning  (?), also significantly contribute to efficiently sampling the configuration space.
@@
Two primary challenges persist in the construction of databases for the development of MLIPs. The first challenge is the comprehensive coverage of defect. For metals, it is relatively straightforward to train MLIPs that can reproduce lattice constants, elastic constants, vacancy formation energy, and energies of simple grain boundaries. However, it remains unclear whether existing MLIPs can effectively model extended defects and their interactions, even if some of them are deemed for general-purpose  (?, ?, ?, ?, ?). This uncertainty primarily stems from the lack of quantification for unknown defects. Although OTF-AL has been proposed to accelerate the development of MLIPs, it is insufficient to sample comprehensive defects in every loading condition. Qi et al. suggested that an ideal strategy would be to efficiently generate and sample configuration space before conducting any DFT calculations  (?). They used M3GNet  (?) as the engine for effective and efficient database generation for atomic hydrogen diffusion in titanium hydrides system. This strategy however is unsuitable for modeling extended defects in metals due to the low accuracy and computational efficiency of universal MLIPs in large-scale simulations.
@@
In this study, we introduce a generalized framework for developing MLIPs specifically designed to model extensive defects in metals. This framework integrates two key components: (1) Empirical Interatomic Potentials-Guided Sampling (EIP-GS), and (2) Periodic Configuration Construction via Grand Canonical Monte Carlo simulations (PCC-GCMC). These components are crucial for overcoming the challenges previously discussed. The EIP-GS method is designed to enhance defect sampling capabilities during large-scale simulations under a wide range of loading conditions. The PCC-GCMC technique, on the other hand, converts non-periodic atomic clusters of selected defects into periodic configurations without the need for vacuum. This conversion is essential for ensuring that these configurations are compatible with standard plane-wave DFT calculations. Focusing on body-centered cubic tungsten (BCC W), our analysis underscores the critical role of the EIP-GS and PCC-GCMC techniques in enhancing the predictive accuracy and reliability of MLIPs for large-scale simulations. Ultimately, we have developed a set of MLIPs optimized for BCC W, capable of simulating a wide range of plastic deformations. We demonstrate the capabilities of our MLIPs through nanoindentation simulations on single crystals. These tests reveal new deformation mechanisms not previously documented, underscoring the robustness and adaptability of our MLIPs in handling complex simulation scenarios.
@@ RESULTS
In this study, we introduce an generalized framework designed to develop versatile MLIPs for accurately modeling extensive defects in metals, as depicted in Fig. 1. This framework is implemented through a sequential three-step process. Initially, a comprehensive pool of defect structures is generated using large-scale simulations based on well-established EIPs, notably the Embedded-Atom Method (EAM) potentials for metals. These potentials have been instrumental for over three decades, enabling the rapid simulation of complex defects across millions of atoms, albeit with lower accuracy compared to DFT. These capabilities are indispensable for simulating a broad spectrum of deformation scenarios, encompassing general GBs in polycrystals, as well as compression, tension, shear, and nanoindentation tests in both polycrystalline and single-crystal systems (Fig. 1a). For a particular metal (BCC W in this study), we employ EAM potentials to generate a diverse set of LAEs for defect, addressing phenomena such as grain boundary relaxation and deformation, dislocation nucleation from material inside and exposed surfaces, dislocation multiplication, and complex dislocation-grain boundary interactions. Subsequently, the D-optimality criterion-base algorithm implemented in MLIP-3  (?) is used to select representative LAEs based on the local atomic neighborhoods. This strategy meticulously evaluates the extensive pool of generated configurations against an established basic dataset (obtained through domain knowledge, as detailed in Supplementary Note 1), enabling the precise selection of the most representative LAEs. This process yields distinct atomic clusters with typical size of 100-200 atoms. For these clusters, vacuum layers of 8 Å thickness are added in all three dimensions of the simulation box to minimize image interaction. We develop an MLIP, named C-MLIP-0, combining the basic dataset with these cluster configurations.
@@
In the second step, our PCC-GCMC strategy is employed to convert all clusters from the first step into periodic configurations as shown in Fig. 1b. For each atomic cluster, we begin by initializing the simulation box size slightly larger than the atomic coordinates in all directions. Subsequently, new atoms (depicted in blue in Fig. 1b) are inserted into the box via GCMC simulations, continuing until no additional atoms can be accommodated. EAM potentials are used in GCMC simulations. During this insertion phase, the core atoms (shown in red) remain fixed to preserve the original LAE of the center atom. After the insertion is complete, both the coordinates of the newly inserted blue atoms and the box size are relaxed simultaneously. This step is crucial to prevent highly distorted LAEs among the newly inserted atoms. This procedure is iteratively repeated, with the box size gradually increased to enable the insertion of atoms into lower energy states. By selecting only the lowest-energy configurations from PCC-GCMC, we ensure that all LAEs in these configurations are physically pertinent to the central LAE. The resulting new DFT dataset, combined with the basic dataset, is utilized to train a foundational model, termed P-MLIP-0, designed to be suitable for simulating common defects.
@@
The final stage, Step 3, is the MLIP refinement phase, as illustrated in Fig. 1c. For specific applications not addressed in the initial EIP-GS of Step 1, such as crack propagation and radiation damage, further EIP-GS and PCC-GCMC operations are required, utilizing appropriate EIPs. A limitation of EIP-GS is its potential bias in generating necessary LAES due to the inherent bias of EIPs. Consequently, new LAEs may be generated during MLIP-based large-scale simulations. To ensure the transferability of MLIPs across diverse applications, we implement the traditional OTF-AL process here for specific cases. All generated atomic clusters are converted into periodic configurations via PCC-GCMC and reincorporated into the training dataset. This process is repeated until no new LAEs are produced by OTF-AL. The final version of our MLIP, named P-MLIP-1, represents the culmination of our development efforts and is designed to robustly model a wide range of defect phenomena.
Significance of PCC-GCMC
In this section, we discuss the significance of PCC-GCMC. To demonstrate the functionality of PCC-GCMC, we employ the example of a screw dislocation core in BCC W, as illustrated in Fig. 2a. The atomic cluster includes a screw dislocation core with three compact core atoms, labeled 1-3. This configuration maintains periodic boundary conditions along the $z$-direction, with free boundary conditions along the $x$ and $y$ directions. The PCC-GCMC process is repeated with varying box dimensions, $\Delta_{x}$ and $\Delta_{y}$, from 1 to 10 Å. Fig. 2b shows the average potential energies of each configuration constructed via PCC-GCMC. Notably, configurations with potential energies below -8.8 eV/atom, depicted in blue, approach the cohesive energy of -8.9 eV/atom which are nearly close to a perfect BCC W. Fig. 2c highlights two representative configurations. In the top panel, we observe a periodic configuration featuring a new non-compact core crossing the periodic boundary along the $y$-axis dislocation, with the remainder of the atoms conforming to the perfect BCC lattice structure, excluding the original screw core. The bottom panel reveals a periodic configuration that includes a screw dislocation dipole, akin to typical screw dislocation dipole configurations with a tilted box. Both configurations depicted in Fig. 2c are generated entirely through automated processes. All newly inserted atoms either form a pure BCC lattice or constitute relevant LAEs, making them suitable for MLIP training.
@@
We generate both the cluster dataset and the periodic dataset following the procedure outlined in Fig. 1a. All large-scale simulations are driven using the EAM-Zhou potential  (?). We then explore the differences between the cluster dataset shown in Fig. 1a and the periodic dataset from Fig. 1b, alongside the basic dataset informed by domain knowledge. This examination is conducted by analyzing the difference in the SOAP descriptor  (?) with parameters set as $r_{\text{cut}}$ = 5 Å, $n_{\text{max}}$ = 12, and $l_{\text{max}}$ = 10  (?). The results are visualized in Fig. 2d, where the first two principal components (PCs) are plotted against each other. In the basic dataset, the majority of atomic environments exhibit PC1 values less than 0, with some data points extending into a long arm, representing surface atoms. The cluster dataset, conversely, divides into two groups: a smaller group that aligns closely with the main portion of the basic dataset and a larger group that positions itself above the extended arm of the basic dataset. This larger group primarily comprises surface atoms, a logical outcome given their proximity to the free surface. The smaller group represents target atoms located at the center of each cluster configuration, which correspond to the LAEs of defects. The size disparity between these groups indicates that only a small fraction of LAEs directly contribute to target defects, while the majority, being surface atoms, do not inform the force and energy calculations of non-surface atoms.
Unlike the cluster dataset, the periodic dataset forms a single cohesive group containing all target atoms from the cluster dataset. There is an overlapping region between the periodic and basic datasets, marked by a dashed line, suggesting shared common LAEs. This overlap signifies that the periodic configurations created via our PCC-GCMC method exhibit characteristics similar to those identified in Ref  (?), with data points in this region likely corresponding to the perfect BCC lattice. Additionally, other data points in the periodic dataset represent inserted atoms from the GCMC process and atoms close to the surfaces in original clusters. Although these LAEs are not explicitly defined in the initial EIP-GS (such as non-compact screw core discovered in Fig. 2c), they may prove instrumental in exploring other types of unknown plastic deformations, which we will discuss further in the following section. Another interesting observation is that the LAEs of defects are not included in the basic dataset (2d), indicating that the existing MLIPs obtained through domain knowledge are unsuitable for modeling general defects.
@@
Next, we proceed to train C-MLIP-0 and P-MLIP-0, utilizing the cluster dataset (C) and the periodic dataset (P), respectively, each integrated with the basic dataset (B). MTP is used for training. The training performance for these models is assessed by comparing the root mean square error (RMSE) of configuration energies, as presented in Fig. 2(e). It is consistently observed that the B+P combination outperforms B+C across various MTP levels. This superior performance of B+P can be attributed to the more straightforward nature of the dataset, which is devoid of the complex, irrelevant LAEs like surface atoms found in the cluster dataset (illustrated in Fig. 2d). This finding underscores the first significant advantage of employing the periodic dataset in MLIP training: the elimination of irrelevant LAEs enhances the training process, leading to more accurate and reliable MLIPs.
Uncertainty quantification across diverse plastic deformations
@@
Our MLIP-0 models, i.e., C-MLIP-0 and P-MLIP-0, are designed to accommodate a wide range of defect simulations including dislocations, GBs, dislocation nucleation and multiplication, and dislocation-GB interactions. To verify the transferability of our MLIP-0 models, we utilize per-atom uncertainty quantification available in MLIP-3  (?) (the extrapolation grade $\gamma$) to evaluate various defects in BCC W. Values of $\gamma$ ranging from 0 to 1 signify interpolation, whereas values exceeding 1 imply extrapolation. Figs 3a-c display three common types of dislocation loops encountered in the plastic deformation of BCC metals: a $\langle 100\rangle$ interstitial loop, a $\langle 111\rangle$ glide loop, and a $\langle 111\rangle$ climb loop. Given that the performance outcomes of the basic dataset combined with cluster one (B+C) and the basic dataset with periodic one (B+P) are comparable, we present only the results for B+C for clarity. Notably, using the basic dataset (B), all atoms within the $\langle 100\rangle$ interstitial loop and those near the edge segments of the $\langle 111\rangle$ glide loop, along with some atoms in the $\langle 111\rangle$ climb loop, exhibit $\gamma$ values significantly greater than 1, indicating a higher level of uncertainty (left panels of Fig. 3a-c). In contrast, all atoms within these dislocation loops, when analyzed using the B+C dataset, show $\gamma$ values less than 1. This demonstrates our MLIP-0 models’ enhanced effectiveness in capturing complex dislocations, with energies and forces of all LAEs being obtainable through interpolation.
@@
Additionally, we assess the transferability of our models regarding general GBs in polycrystals and their temperature-induced evolution. We calculate the extrapolation grade $\gamma$ for all the atoms in a polycrystal relaxed by a MEAM potential  (?) at 300 K and 2000 K. It should be noted that our EIP-GS in the initial EIP-GS only use EAM-Zhou potential  (?), such that the GBs obtained by MEAM may have different LAEs. Fig. 3d illustrates the polycrystal at 300 K, with each atom color-coded according to $\gamma$ values obtained from B+C. Notably, all atoms in this representation exhibit $\gamma$ values less than 1, indicating low uncertainty and high confidence in the predictions. Figs 3e and f display histogram plots of the $\gamma$ values for this polycrystal calculated using different datasets: the basic dataset (B), a dataset from a Gaussian Approximation Potential (GAP, see the Supplementary Note 1 for the details)  (?), B+C, and B+P at 300 K and 2000 K. The analysis reveals that both datasets B and GAP predict $\gamma$ values greater than 2 for atoms at both 300 K and 2000 K, indicating higher uncertainty. This is particularly pronounced for the dataset B. In contrast, $\gamma$ values calculated by datasets B+C and B+P are consistently below 1 at 300 K. However, at 2000 K, there are a few exceptions with values of $\gamma$ slightly higher than 2. This indicates that although the models perform well at lower temperatures, at higher temperatures the uncertainty escalates but stays manageable. These results underscore the high transferability and effectiveness of generated defect genome in simulating the behavior of general GB and the thermodynamics of polycrystals.